OpenStudy (anonymous):

zy2-9y+3=0 can someone help me solve this im getting stuck where i plug in at -9^2-4(1)(3)

OpenStudy (anonymous):

\[x = \frac{ 9 \pm \sqrt{(-9)^{2}-4(1)(3)} }{ 2(1)} = \frac{\sqrt{ 81-12} }{ 2} \]

OpenStudy (anonymous):

ok i worked it out and got 9+- 8 /2

OpenStudy (anonymous):

You should get: \[x = \frac{ 9\pm \sqrt{69} }{ 2 }\] This is as far as it simplifies unless you are asked to find an approximate value.